A train leaves Philadelphia at 3:00pm. A sexond train leaves the same city in the same direction at 9:00pm. the second train travels 120 mph faster than the first. If the second train overtakes the first at 11:00pm, what is the speed of each of the two trains?
Answer #1

Let speed of fast train is x mph

Speed of second train is x + 120 mph

Fast train travels (11 - 3) = 8 hour

2nd train travel (11 - 9) = 2 hour

According to the question

Distance travelled by first train in 8 hour = distance travelled by second into hour.

so, 8x = 2(x+120)

or, 8x = 2x +240

or, 8x - 2x = 240

or, 6x = 240

or, x = 40 [divide both side by 6]

Speed of fast train is 40 mph

Speed of second train is (40 + 120 )=160 mph